# HiVG: Hierarchical Multimodal Fine-grained Modulation for Visual Grounding

### v1

@@ 4.2. Comparison with State-of-the-Art Methods
Experimental Setting. It is worth emphasizing that, as described in Sec. 2.1, our focus is on the transfer learning of self-supervised pre-trained models for grounding tasks. We follow the basic fine-tuning setting with the same as CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), etc.. In particular, we also contrast with the traditional setting of fine-tuning with pre-trained detection and language models (e.g., TransVG (Deng et al., 2021), TransVG++(Deng et al., 2023), etc.). Additionally, we also follow the previous works that utilized a mixed dataset pre-training setting (e.g., MDETR (Kamath et al., 2021), OFA(Wang et al., 2022b)) and mix the training data (only includes the RefCOCO/+/g, ReferIt, Flickr30k datasets) for intermediate pre-training. This allows us to compare our results with these works in a relatively fair manner. The details are presented in Tab. 1.
@@ 4.3. Ablation Study
Ablation Study of MACB. As shown in Tab. 4, we conducted an ablation study on the implementation of the multi-layer adaptive cross-modal bridge (MACB, default using 12 layers of text features). The weights in the table denotes the sample-agnostic weights. The table shows that our designed structure can effectively utilize multi-level text features and achieve hierarchical adaptation. More detailed results and descriptions will be provided in the Appendix.
Ablation Study of Hi LoRA. As presented in Tab. 5 and Fig. 4-(d), we conducted an ablation study on Hi LoRA with different LoRA stages and various low ranks. It is observed that employing 3-stage Hi LoRA with low rank as 32 achieves the best performance. More detailed results will be provided in the Appendix.
[…]
Architecture of BeiT-3 (384$\times$384) val test
two-tower w. frozen 38.03 38.38
two-tower w. vanilla LoRA 64.36 64.73
two-tower w. Hi LoRA 67.11 67.39
one-tower w. frozen 44.83 46.03
one-tower w. vanilla LoRA 74.65 75.32
one-tower w. Hi LoRA 77.19 78.30
Table 7. Generality study of Hi LoRA in object detection (b) and instance segmentation (m) on MS-COCO dataset.
[…]
vanilla CLIP w. backbone frozen 640 33.4 52.6 27.7 48.2
CLIP w. vanilla LoRA 640 44.1 63.2 38.5 60.7
CLIP w. Hi LoRA 640 47.5 67.1 43.1 64.6
MaskCLIP (Dong et al., 2023) (baseline, CVPR’23) 1333 45.4 64.7 40.9 62.1
4.4. Discussion on the Generality of Hi LoRA
[…]
@@ 4.5. Qualitative Results
We visually present the results of several relatively challenging examples in Fig. 5. The attentions show the [REG] token over vision tokens from the last grounding block of each model. HiVG demonstrates exceptional semantic understanding capabilities in the complex sentences. More examples are provided in Appendix.
5. Conclusion and Future Work
@@
Table 8. The detailed statistics of RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), RefCOCOg (Mao et al., 2016), ReferItGame (Kazemzadeh et al., 2014b) and Flickr30K Entities (Plummer et al., 2015) datasets. We represent test split and testA split in the same column.
@@ Appendix C Extra Result Analysis
Analysis of the Computational Complexity in Figure 4-(b) of the Main Text. According to Table 2 of the main text, the number of parameters in existing models (except for QRNet (Ye et al., 2022a)) is not significantly different, roughly ranging from 150M to 210M. However, the computational complexity of the Transformer architecture heavily depends on the length of input token sequences, i.e., there is an $\mathcal{O}(n^{2})$ complexity. For example, TransVG++ (Deng et al., 2023) utilizes a 640$\times$640 resolution image as input with a patch size of 16$\times$16, resulting in a sequence length of $(640/16)^{2}=40^{2}=1600$ in the vision backbone. In contrast, our HiVG employs a smaller resolution image of 224$\times$224 with a patch size of 16$\times$16; thus, our visual sequence length is only $(224/16)^{2}=14^{2}=196$. Taking [CLS] and [REG] tokens into account, HiVG’s vision sequence length is merely $(196+1)/(1600+2)=12.29\%$ compared to that of TransVG++ (i.e., TransVG++ is 8.13×8.13$\times$8.13 × larger than HiVG), demonstrating a dominant difference. Unlike the other works (Deng et al., 2023; Su et al., 2023), our framework can obtain state-of-the-art results without relying on high-resolution images. This significantly reduces the calculation complexity and greatly accelerates the training and reasoning computation of our HiVG framework.

### v2

@@ 4.2. Comparison with State-of-the-Art Methods
Experimental Setting. It is worth emphasizing that, as described in Sec. 2.1, our focus is on the transfer learning of self-supervised pre-trained models for grounding tasks. (1) We follow the basic fine-tuning setting with the same as CLIP-VG (Xiao et al., 2023) and Dynamic-MDETR (Shi et al., 2022), etc.. (2) In particular, we also compare with the traditional setting of fine-tuning with pre-trained detection models (e.g., TransVG (Deng et al., 2021), TransVG++(Deng et al., 2023), etc.). (3) Additionally, we also follow the previous works that utilized a dataset-mixed pre-training setting (e.g., MDETR (Kamath et al., 2021), OFA(Wang et al., 2022b)) and mix the training data (only includes the RefCOCO/+/g, ReferIt, Flickr30k datasets) for intermediate pre-training. This allows us to compare our results with these works in a relatively fair manner. The details are presented in Tab. 1.
@@ 4.3. Ablation Study
Ablation Study of MACB. As shown in Tab. 4, we conducted an ablation study on the implementation of the multi-layer adaptive cross-modal bridge (MACB, default using 12 layers of text features). The weights in the table denotes the sample-agnostic weights. The table shows that our designed structure can effectively utilize multi-level text features and achieve hierarchical adaptation.
Ablation Study of HiLoRA. As presented in Tab. 5 and Fig. 4-(d), we conducted an ablation study on HiLoRA with different LoRA stages and various low ranks. It is observed that employing 3-stage HiLoRA with low rank as 32 achieves the best performance.
@@ 4.4. Qualitative Results
We visually present the results of several relatively challenging examples in Fig. 5. The attentions show the [REG] token over vision tokens from the last grounding block of each model. HiVG demonstrates exceptional semantic understanding capabilities in the complex sentences.
5. Conclusion
@@
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grants 62036012, U23A20387, 62322212, 62072455, in part by Pengcheng Laboratory Research Project under Grant PCL2023A08, and also in part by National Science and Technology Major Project under Grant 2021ZD0112200.
@@ Acknowledgements.
Table 6. The detailed statistics of RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), RefCOCOg (Mao et al., 2016), ReferItGame (Kazemzadeh et al., 2014b) and Flickr30k(Plummer et al., 2015) datasets. We represent test and testA split in same column.
@@ Appendix B Implementation Details
Appendix C Extra Result Analysis
@@
Analysis of the Computational Complexity in Figure 4-(b) of the Main Text. According to Table 2 of the main text, the number of parameters in existing models (except for QRNet (Ye et al., 2022a)) is not significantly different, roughly ranging from 150M to 210M. However, the computational complexity of the Transformer architecture heavily depends on the length of input token sequences, i.e., there is an $\mathscr{O}(n^{2})$ complexity. For example, TransVG++ (Deng et al., 2023) utilizes a 640$\times$640 resolution image as input with a patch size of 16$\times$16, resulting in a sequence length of $(640/16)^{2}=40^{2}=1600$ in the vision backbone. In contrast, our HiVG employs a smaller resolution image of 224$\times$224 with a patch size of 16$\times$16; thus, our visual sequence length is only $(224/16)^{2}=14^{2}=196$. Taking [CLS] and [REG] tokens into account, HiVG’s vision sequence length is merely $(196+1)/(1600+2)=12.29\%$ compared to that of TransVG++ (i.e., TransVG++ is 8.13×8.13$\times$8.13 × larger than HiVG), demonstrating a dominant difference. Unlike the other works (Deng et al., 2023; Su et al., 2023), our framework can obtain state-of-the-art results without relying on high-resolution images. This significantly reduces the calculation complexity and greatly accelerates the training and reasoning computation of our HiVG framework.